# Neural Diversity Regularizes Hallucinations in Small Language Models

### v1

@@ 4.4 Task-Dependent Optimality
Table 2: The optimal amount of neural diversity varies by task. Hallucination tasks show large gains with optimal $P\in\{2,4,8\}$ (up to +25.6%), while knowledge tasks mostly peak at $P=1$. Even among hallucination tasks, optimal $P$ differs: HaluEval variants peak at $P=4$, TruthfulQA at $P=2$, MemoTrap at $P=8$. Each row shows the performance peak (best $P$), best score, relative increase vs. $P=1$ baseline, and statistical significance (*** for $p<0.001$, * for $p<0.05$).
@@ 5 Mechanistic Analysis
5.1 Neural Diversity as the Causal Mediator
Experiment Design. Starting with a pre-trained ND-LoRA $P=4$ model, we inject a corruption hook at the RMSNorm layer that randomly substitutes the hidden state at randomly-chosen positions in a given stream from another stream, perturbing $\mathcal{D}_{\text{spec}}$ while preserving activation magnitudes. We evaluate on a matched basis: each corrupted evaluation is paired with an uncorrupted baseline using identical samples and resampling indices. Across 4 sub-experiments with different random seeds, we collect N=128 paired samples per task. This paired design maximizes statistical power by controlling sample-level variance, analyzed via paired t-tests with Fisher meta-analysis.
Results. Table 3 provides statistically robust evidence that neural diversity causally affects performance. All three tasks show highly significant accuracy drops ($p<0.001$) when stream-level substitution perturbs diversity ($\Delta\mathcal{D}_{\text{spec}}\approx 0.025$). While effect sizes are modest (0.3% to 0.7% score reduction) — likely because artificial stream substitution creates out-of-distribution corruption patterns — the statistical significance establishes causality beyond correlational association.
Task $\Delta\mathcal{D}_{\text{spec}}$ $\Delta$ Score SE d p-value Sig. N
@@ 5.1 Neural Diversity as the Causal Mediator
Table 3: Artificial corruption of neural diversity establishes statistical causality. Perturbing neural diversity ($\Delta\mathcal{D}_{\text{spec}}>0$) causes accuracy drops across tasks with high statistical significance ($p<0.001$) via paired t-tests with Fisher meta-analysis (N=4 sub-experiments × 128 samples each).
@@ 5.2 Ablations
To isolate the contributions of ND-LoRA, we systematically ablate ND-LoRA components at fixed $P=4$ streams. All variants maintain parameter parity through LoRA rank adjustments, enabling fair comparison. We measure run-time spectral diversity ($\mathcal{D}_{\text{spec}}$) at the aggregation layer using evaluation samples, quantifying actual cross-stream correlation during inference.
@@ 5.3 Computational Considerations
Variant Streams LoRA Regul. Target $\mathcal{D}_{\text{spec}}$ $\overline{\Delta}$% Score $\Delta$ Cost
@@
Stream LoRA $P$ Stream D All 0.2678 +2.9% 1.00003x / 1.1x
Stream LoRA-BT $P$ Stream D + BT All 0.1400 +4.9% 1.00004x / 1.1x
ND-LoRA $P$ Stream D + BT KVQ 0.3755 +12.8% 1.00004x / 1.1x
Table 4: Ablations reveal super-linear combination of impact. Stream LoRA (+2.9%) and Barlow Twins (+1.4%) combine super-linearly (+4.9%), and focusing on KVQ attention amplifies to +12.8%. LoRA: single shared vs. $P$ stream-aware adapters. Regularization: Dropout vs. Barlow Twins. Target: All layers vs. KVQ attention only. $\mathcal{D}_{\text{spec}}$: spectral diversity (lower better). $\overline{\Delta}$% Score: avg. change (hallucination benchmarks).

### v2

# Neural Diversity Regularizes Hallucinations in Language Models

@@ 1 Introduction
Table 1: Optimal neural diversity is task-dependent: hallucination tasks benefit from neural diversity, knowledge tasks do not. De-aggregating Figure 1, hallucination benchmarks consistently show large gains with increased diversity (up to 25.6%, HaluEval-Summ, $P_{\star}=4$), while knowledge retrieval mostly peaks at $P_{\star}=1$. This asymmetry supports hallucination as a reliability problem distinct from factual recall. Significance: *** $p<0.001$, * $p<0.05$.
Despite scaling to trillions of parameters, language models hallucinate persistently (Lin et al., 2021). This reliability crisis is acute for small language models — increasingly favored for edge and agentic use cases (Zheng et al., 2025; Belcak et al., 2025) — whose compressed representations make them especially vulnerable to hallucinations, with even well-resourced efforts like GPT-OSS 20B exhibiting 91% hallucination rates on factual benchmarks (OpenAI, 2025).
Current hallucination mitigation strategies are largely empirically driven but theoretically ungrounded and target average performance rather than tail risk. RLHF optimizes mean harmlessness (Bai et al., 2022), RAG improves average factual grounding (Niu et al., 2024), and contrastive decoding enhances mean generation quality (Li et al., 2023b). While inference-time approaches like self-consistency and LoRA ensembling (Wang et al., 2022; 2023) reduce hallucinations through diverse sampling, they lack formal tail-probability guarantees. Similarly, parallel scaling methods (Chen et al., 2025) target first-moment improvements in perplexity and task accuracy. Yet controlling catastrophic failures requires bounding the tails of $\mathbb{P}(\text{hallucination})$, not just optimizing mean behavior.
Formal ensemble theory exists but targets the wrong objective. Classical ensemble methods (Krogh & Vedelsby, 1994) provide rigorous diversity theory to reduce mean generalization error $\operatorname*{\mathbb{E}}[\text{loss}]$, not tail-probability bounds for hallucinations. Deep ensembles (Lakshminarayanan et al., 2017) quantify uncertainty but lack hallucination-specific guarantees. Without explicit diversification, parallel architectures suffer representational collapse (Jing et al., 2022), leaving reliability gains unrealized.
To our knowledge, we provide the first formal framework for hallucination probability tail bounds in ensembled language models, reframing it as a second-moment reliability problem. Drawing on portfolio theory (Markowitz, 1952), we prove that decorrelated parallel representations (neural diversity) reduce this tail bound and introduce ND-LoRA (Neural Diversity Low-Rank Adaptation) to concretely demonstrate its hallucination reduction capabilities.
Our contributions are:
  • Theoretical Linkage: We reframe hallucinations as a second-moment reliability problem and prove (i) a portfolio-theoretic bound showing hallucination probability $\mathbb{P}(\mathrm{H})~\propto~1/P$ with $P$ decorrelated parallel representations (Theorem 1); and, (ii) non-monotonicity in reliability scaling (Theorem 2), showing that excessive parallelism can degrade diversity (and thus reliability) under common circumstances. We further show (iii) our theoretical predictions achieve $R^{2}=0.962$ in fitting empirical reliability gains (Figure 1), establishing quantitative validation rare in neural hallucination research.
[…]
  • Mechanistic Analysis: We establish that neural diversity mediates hallucination in four ways: (i) causality via perturbation ($p<0.001$, Table 3), (ii) quantitative scale via correlation (+0.1% diversity $\Leftrightarrow$ -3.8% hallucination, Figure 3), (iii) super-linear effects via ablation (Table 4), and (iv) task-dependent optima via scaling sweeps (Table 1).
2 A Theory of Neural Diversity
Why don’t existing scaling methods improve reliability? Without explicit diversity mechanisms, gradient descent drives parallel streams toward similar representations through representational collapse (Jing et al., 2022), leaving reliability gains unrealized. We establish the first hallucination tail bounds for ensembled language models, proving that neural diversity reduces hallucinations and providing mathematical foundations for ND-LoRA.
Our strategy adapts portfolio theory to neural architecture design. Classical ensemble methods reduce mean error $\operatorname*{\mathbb{E}}[\text{loss}]$ through variance reduction (Krogh & Vedelsby, 1994), treating correlation as a factor that limits accuracy gains. In contrast, portfolio theory manages tail risk — rare but catastrophic failures — by diversifying across correlated assets (Markowitz, 1952). We adapt the latter framework to tail bound hallucination probability $\mathbb{P}(\text{hallucination})$, where correlation becomes the primary control variable for reliability rather than a secondary constraint on mean performance.
2.1 Preliminaries
Modern language models hallucinate by fabricating facts, generating content inconsistent with input, or creating unsupported claims (Maynez et al., 2020; Ji et al., 2023). While comprehensive taxonomies exist (Huang et al., 2024), we model hallucinations through a simple signal-noise proxy that captures the underlying reliability failure (Figure 1) while remaining analytically tractable.
Signal-noise model.
Let $x\in X$ be a query with oracle output $y_{\star}(x)\in\mathbb{R}^{V}$ and corresponding hidden representation $z_{\star}(x)\in\mathbb{R}^{d}$. Consider an architecture that employs $P$ parallel computational pathways called streams, each processing the same input $X$ through the same model but in perturbed ways. We model the hidden output of each stream as $Z_{i}=z_{\star}+\varepsilon_{i}$ where $\varepsilon_{i}\in\mathbb{R}^{d}$ is centered noise with variance $\sigma_{i}^{2}>0$. The noise covariance $\Sigma\in\mathbb{R}^{P\times P}$ has entries $\Sigma_{ij}\triangleq\operatorname*{\mathbb{E}}[\langle\varepsilon_{i},\varepsilon_{j}\rangle]$ with pairwise correlations $\rho_{ij}\triangleq\Sigma_{ij}/(\sigma_{i}\sigma_{j})$ for $i\neq j$. We aggregate hidden representations via $\widehat{Z}_{w}=\sum_{i}w_{i}Z_{i}$ with weights summing to one. For readability, we omit $x$ where obvious and denote the average noise variance by $\bar{\sigma}^{2}\triangleq\operatorname*{\mathbb{E}}[\sigma_{i}^{2}]$ and average correlation $\bar{\rho}\triangleq\operatorname*{\mathbb{E}}_{i<j}[\rho_{ij}]$.
High-dimensional structure.
High-dimensional representations exhibit predictable geometric regularity that we exploit for analysis. We assume: (i) Lipschitz decoding, where outputs $\widehat{Y}_{w}(x)=f(\widehat{Z}_{w}(x))$ and $y_{\star}(x)=f(z_{\star}(x))$ satisfy $\|f(z)-f(z^{\prime})\|_{2}\leq L\|z-z^{\prime}\|_{2}$ for some $L>0$; and, (ii) norm concentration, where $\|\tilde{z}_{i}(X)\|^{2}_{2}\approx d$ with small relative variance for per-feature whitened representations $\tilde{z}_{i}$. Both properties are standard in high-dimensional probability (Vershynin, 2018) and neural network analysis (Fazlyab et al., 2019; Bartlett et al., 2017).
Neural representations.
At a chosen design layer, each stream exposes a $d$-dimensional representation $z_{i}(X)$. We whiten per-feature to obtain $\tilde{z}_{i}$ with zero mean and identity covariance. For streams $i<j$, the cross-correlation matrix is $C^{(ij)}\triangleq\operatorname*{\mathbb{E}}\big[\tilde{z}_{i}\tilde{z}_{j}^{\top}\big]\in\mathbb{R}^{d\times d}$ whose diagonal entries measure same-feature similarity and off-diagonal entries capture cross-feature alignment. Finally, using the widely-exploited observation that trained networks exhibit locally linear behavior at their operating point (Goodfellow et al., 2015; Simonyan et al., 2014), we connect representations to noise via local linearity: $\xi_{i}\;=\;A\tilde{z}_{i}$ for a shared linear readout $A\in\mathbb{R}^{d\times d}$ with finite condition number $\kappa$.
Neural diversity index.
We define a simple cosine-based index to measure cross-stream diversity:
$\mathcal{D}\triangleq\sqrt{\operatorname*{\mathbb{E}}_{i<j}\left[\frac{(\tilde{z}_{i}\cdot\tilde{z}_{j})^{2}}{\|\tilde{z}_{i}\|^{2}\|\tilde{z}_{j}\|^{2}}\right]}.$ (1)
Lower $\mathcal{D}$ indicates greater neural diversity: $\mathcal{D}=0$ means all streams are perfectly orthogonal, while $\mathcal{D}=1$ means streams have suffered complete collapse.
Hallucinations.
We define the output error as $E_{w}\triangleq\|\widehat{Y}_{w}(x)-y_{\star}(x)\|_{F}$, which is comparable to metrics like TruthfulQA-MC2 (Lin et al., 2021). For tolerance $\delta>0$, the hallucination event is $\mathrm{H}_{\delta}\triangleq\{E_{w}\geq\delta\}$. Our goal is to bound $\mathbb{P}(\mathrm{H}_{\delta})$ as a function of neural diversity $\mathcal{D}$ across streams $P$.
2.2 Neural Diversity Bounds Hallucination
Classical portfolio theory (Markowitz, 1952) gives the variance of an equally weighted portfolio of $P$ assets with average variance $\bar{\sigma}^{2}$ and average pairwise correlation $\bar{\rho}$ as:
$\mathrm{Var}(Y)\;=\;\bar{\sigma}^{2}\!\left(\frac{1-\bar{\rho}}{P}+\bar{\rho}\right).$ (2)
To use this observation for hallucinations, we must first connect neuron-level representations to portfolio-level noise correlations. Exploiting the fact that (i) our ensemble has one underlying model with aligned neuron-level representations and (ii) our model has geometric regularity in representation and output, the following lemma establishes this mapping:
Lemma 1 (Average Correlation Bound).
Suppose there exists a constant $C_{4}\geq 1$ such that $\operatorname*{\mathbb{E}}[\|\xi_{i}\|_{2}^{4}]\leq C_{4}\,\sigma_{i}^{4}$ for all $i$. Then the average pairwise noise correlation satisfies
$|\bar{\rho}|\;\leq\;C_{\ast}\,\mathcal{D},$ (3)
where $C_{\ast}\triangleq\sqrt{C_{4}}\,\kappa^{2}$ depends only on the kurtosis bound and the readout condition number $\kappa$.
Proof sketch.
We proceed in two steps: (1) Spectral bounds imply linear readout distorts cosines by at most $\kappa^{2}$, so noise-space diversity $\mathcal{D}_{\xi}\leq\kappa^{2}\mathcal{D}$. (2) Cauchy–Schwarz twice — inner products to cosines, then kurtosis — gives $|\rho_{ij}|\leq\sqrt{C_{4}}\,\mathcal{D}_{\xi,ij}$; averaging pairs completes the proof. ∎
We now have a direct path to tail-bound $P(\mathrm{H}_{\delta})$ as a function of $\mathcal{D}$ and $P$. For readability, we assume uniform weights $w_{i}=1/P$ below but our approach can also be easily applied to arbitrary weights.
Theorem 1 (Hallucination Bound with Diversity).
For any tolerance $\delta>0$,
$\mathbb{P}(\mathrm{H}_{\delta})\;\leq\;\frac{\frac{1-C_{\ast}\,\mathcal{D}}{P}+C_{\ast}\,\mathcal{D}}{\frac{1-C_{\ast}\,\mathcal{D}}{P}+C_{\ast}\,\mathcal{D}+SNR},$ (4)
where $SNR\triangleq\delta^{2}/\bar{\sigma}^{2}$.
Proof sketch.
Lemma 1 bounds $|\bar{\rho}|\leq C_{\ast}\,\mathcal{D}$, linking noise variance to representational diversity. Plugging into Equation 2, applying Chebyshev and normalizing by $\bar{\sigma}^{2}$ yields the stated bound. ∎
This completes the first half of our theoretical result: Neural diversity mediates hallucination probability. With perfect de-correlation ($\bar{\rho}=0$), hallucination probability scales as $O(1/P)$ — more streams reduce hallucination risk. When streams collapse ($\bar{\rho}=1$), the bound becomes independent of $P$, explaining why naive ensembling without diversification provides no reliability benefits.
2.3 Non-Monotonic Scaling Behavior
Theorem 2 (U-shaped Behavior).
Suppose $\bar{\rho}(P)=\rho_{0}+\beta(P-1)^{\gamma}$ for constants $\rho_{0}\in[0,1)$, $\beta>0$, $\gamma>0$. Define
$v(P)\triangleq\mathrm{Var}(E_{w})=\bar{\sigma}^{2}\left(\frac{1-\bar{\rho}(P)}{P}+\bar{\rho}(P)\right),\qquad B(P)\triangleq\frac{v(P)}{v(P)+\delta^{2}}.$ (5)
Then $B(P)$ is U-shaped: there exists $P_{\star}\geq 1$ minimizing $\mathbb{P}(\mathrm{H}_{\delta})$, with $P_{\star}$ controlled by how fast $\bar{\rho}(P)$ degrades with $P$.
Proof sketch.
The hallucination bound $B(P)$ is monotonic in variance $v(P)$, so we analyze $v(P)$ directly. There are two competing effects: the $1/P$ term drives variance down, while growing correlation $\bar{\rho}(P)=\rho_{0}+\beta(P-1)^{\gamma}$ eventually dominates. Differentiating shows $v^{\prime}(P)$ changes sign exactly once, yielding a unique minimum $P_{\star}$ whose location depends $\beta,\gamma\text{ and }\rho_{0}$. ∎
This theorem establishes non-monotonicity — hallucination probability $\mathbb{P}(\mathrm{H}_{\delta})$ actually increases for larger $P$, meaning reliability degrades. This is stronger than the well-known diminishing returns of ensembles (where improvement slows but continues). While ensemble theory also shows optimal size matches the number of class labels for accuracy-optimized classifiers (Bonab & Can, 2019), we prove and validate (Figure 1) that diversity can degrade in generative language models with excessive parallelism under common circumstances and also harm reliability.
2.4 Theoretical Validation
By measuring empirical diversity $\mathcal{D}(P)$ and plugging these values into Theorem 1’s bound, we achieve $R^{2}=0.962$ (Figure 1), explaining 94.3% of empirical reliability variation. This fit uses only two free parameters ($C_{\ast}$, SNRSNR) shared across all tasks and observations, with $\mathcal{D}(P)$ fixed from empirical measurements. Theorem 2’s correlation growth model provides a mechanism for observed concavity: correlation grows as $O((P-1)^{\gamma})$, overwhelming the $O(1/P)$ diversification benefit. This alignment — rare in hallucination research where theory often lags empirics — validates our portfolio-theoretic framework.
Together, Theorem 1 and Theorem 2 show that (i) reducing $\mathcal{D}$ reduces hallucinations and (ii) there exists an optimal $P_{\star}$ that minimizes hallucinations. Next, we show how to construct an architecture and training protocol to reduce $\mathcal{D}$ and find $P_{\star}$.
3 ND-LoRA: A Practical Demonstration
We introduce ND-LoRA (Neural Diversity Low-Rank Adaptation), a parameter-efficient method that demonstrates our theoretical framework for neural diversity regularization. ND-LoRA extends the ParScale architecture with stream-aware LoRA adapters and explicit decorrelation objectives. Figure 2 visually summarizes our approach.
3.1 Architecture
$y=\text{LM\_Head}\left(\sum_{i=1}^{P}w_{i}\cdot h_{i}^{(L)}\right)$ (7)
where $w_{i}=(1-\varepsilon)\cdot\text{softmax}(\text{MLP}([h_{1}^{(L)},\ldots,h_{P}^{(L)}]))_{i}+\varepsilon/P$ are dynamic weights with label smoothing ($\varepsilon=0.1$) computed from the concatenated stream representations. This prevents attention collapse by ensuring minimum weight $\varepsilon/P$ for each stream.
3.2 Barlow Twins Regularization
To encourage neural diversity, we apply Barlow Twins regularization across all pairs of streams $i<j$ at a pre-specified design layer $\ell_{\star}$.
Let $z_{i}\in\mathbb{R}^{B\times T\times d}$ denote the hidden representations of stream $i$ at the design layer for a batch of size $B$ and sequence length $T$. We first apply batch normalization and mean-centering to obtain whitened features $\tilde{z}_{i}$. We then calculate the cross-correlation matrices $C^{(ij)}\in\mathbb{R}^{d\times d}$ as in subsection 2.2 and apply standard Barlow Twins (Zbontar et al., 2021) for each pair of streams $i<j$:
$\mathcal{L}_{BT}=\mathop{\operatorname*{\mathbb{E}}}_{i<j}\left\lVert C^{(ij)}-I\right\rVert_{F}$ (8)
$\mathcal{L}=\mathcal{L}_{CE}+\lambda_{BT}\mathcal{L}_{BT}$ (9)
4 Experimental Validation
We validate ND-LoRA through systematic hallucination reduction experiments using parameter- and data-matched comparisons. We describe our full experimental setup in subsection A.5.
4.1 Key Results
Model HaluEval MemoTrap TruthfulQA NQ Wikitext WG
ND-LoRA R16 (P=2) 0.481* 0.666* 0.442* 0.055 0.784 0.574
ParScale R32 (P=2) 0.439 0.638 0.412 0.059 0.793 0.564
Qwen LoRA R32 0.400 0.634 0.403 0.065 0.778 0.572
Table 2: Even at $P=2$ streams, ND-LoRA achieves up to 20.2% relative hallucination reduction vs. parameter-matched baseline. Across hallucination benchmarks, ND-LoRA shows statistically significant improvements (HaluEval-Summarization, MemoTrap, TruthfulQA-MC2) while maintaining competitive Winogrande, NQ, and Wikitext BPB (lower is better) general-purpose capabilities. Baselines use higher LoRA ranks for parameter parity. * indicates $p<0.05$.
Although ND-LoRA’s improvements specifically target reliability benchmarks, they preserve general capabilities. Qwen slightly outperforms on Wikitext (0.778 vs. 0.784) and Natural Questions (0.065 vs. 0.055), but ND-LoRA wins slightly on Winogrande (0.574 vs. 0.572).
Figure 3: Reliability improves as neural diversity increases (lower $\mathcal{D}$). Specifically, diversity ($\mathcal{D}$) is negatively correlated with HaluEval-Summarization performance (slope=-37.842, R²=0.237, p=0.002), consistent with $\mathbb{P}(\text{H})\propto\mathcal{D}$ in Theorem 1.
These findings establish neural diversity as a practical reliability mechanism. Consistent improvements across hallucination benchmarks with preserved general performance suggest ND-LoRA addresses fundamental reliability challenges rather than metric-specific optimization. Figure 3 demonstrates strong empirical correlation between neural diversity and performance, building intuition for the causal relationship established in subsection 5.1.
4.2 Task-Dependent Optimality
@@ 5 Mechanistic Analysis
Task $\Delta\mathcal{D}$ $\Delta$ Score SE d p-value Sig. N
@@
Table 3: Artificial corruption of neural diversity establishes statistical causality. Perturbing neural diversity ($\Delta\mathcal{D}>0$) causes accuracy drops across tasks with high statistical significance ($p<0.001$) via paired t-tests with Fisher meta-analysis (N=4 sub-experiments × 128 samples each).
5.1 Neural Diversity as the Causal Mediator
Experiment Design. Starting with a pre-trained ND-LoRA $P=4$ model, we inject a corruption hook at the RMSNorm layer that randomly substitutes the hidden state at randomly-chosen positions in a given stream from another stream, perturbing $\mathcal{D}$ while preserving activation magnitudes. We evaluate on a matched basis: each corrupted evaluation is paired with an uncorrupted baseline using identical samples and resampling indices. Across 4 sub-experiments with different random seeds, we collect $N=128$ paired samples per task. This paired design maximizes statistical power by controlling sample-level variance, analyzed via paired t-tests with Fisher meta-analysis.
Results. Table 3 provides statistically robust evidence that neural diversity causally affects performance. All three tasks show highly significant accuracy drops ($p<0.001$) when stream-level substitution perturbs diversity ($\Delta\mathcal{D}\approx 0.025$). While effect sizes are modest (0.3% to 0.7% score reduction) — likely because artificial stream substitution creates out-of-distribution corruption patterns — the statistical significance establishes causality beyond correlational association.
@@ 5.2 Ablations
To isolate the contributions of ND-LoRA, we systematically ablate ND-LoRA components at fixed $P=4$ streams. All variants maintain parameter parity through LoRA rank adjustments, enabling fair comparison. We measure inference-time diversity ($\mathcal{D}$) at the aggregation layer using evaluation samples, quantifying actual cross-stream correlation during inference.
@@
Notably, ParScale’s original work found prefix tuning superior to LoRA for mean loss (Table 6 in Chen et al. 2025). However, stream-aware LoRA is necessary for reducing tail probability: even with Barlow Twins, prefix tuning collapses streams ($\mathcal{D}=0.9988$), while stream-aware LoRA enables decorrelation ($\mathcal{D}=0.1530$). This illustrates how second-moment objectives require different architectural choices than first-moment objectives.
@@ 5.3 Computational Considerations
5.4 Practical Approximability
While task-optimal $P_{\star}$ varies (Table 1), practitioners need not search exhaustively. Defaulting to $P=4$ achieves 96% of oracle performance across all benchmarks. Additionally, a simple router (subsection A.7) achieves 97% by predicting $P$ from prompt statistics, revealing a retrieval-vs-verifiability tradeoff: question-dense prompts favor low $P$, while longer prompts favor higher $P$.
Variant Streams LoRA Regul. Target $\mathcal{D}$ $\overline{\Delta}$% Score $\Delta$ Cost
@@ 5.4 Practical Approximability
Stream LoRA $P$ Stream D All 0.3544 +2.9% 1.00003x / 1.1x
Stream LoRA-BT $P$ Stream D + BT All 0.1530 +4.9% 1.00004x / 1.1x
ND-LoRA $P$ Stream D + BT KVQ 0.4112 +12.8% 1.00004x / 1.1x
Table 4: Ablations reveal super-linear combination of impact. Stream LoRA (+2.9%) and Barlow Twins (+1.4%) combine super-linearly (+4.9%), and focusing on KVQ attention amplifies to +12.8%. LoRA: single shared vs. $P$ stream-aware adapters. Regularization: Dropout vs. Barlow Twins. Target: All layers vs. KVQ attention only. $\mathcal{D}$: Neural Diversity Index (lower is better). $\overline{\Delta}$% Score: avg. change (hallucination benchmarks). Ablations shown at fixed $P=4$ streams.